# A General Model for Static Contact Angles

### v1

@@ 1 Introduction
Prior work has focused on developing unified models that bridge the pure Wenzel and Cassie-Baxter extremes and extend these classical models.46; 47; 48 These efforts employ free energy minimization to account for multiple wetting states and assess their thermodynamic stability, for specific wetting configurations and surface geometries such as a macroscale sessile droplets on much smaller micropillar arrays. Despite the significant progress on the fundamental and applied aspects, open questions remain, namely how to predict static contact angles and the hysteresis range from a knowledge of topographic parameters, surface chemistry aging, and ambient conditions affecting the degree of liquid infiltration on the surface below and ahead of the wetting liquid phase. In an effort to address this, this work proposes a general and simple-to-use model, expanding Young’s local thermodynamic approach to consider surfaces with topographic features of general geometry and varying degrees of liquid infiltration. The introduced unified model recovers classical wetting limits (Wenzel, Cassie-Baxter, and hemiwicking states), accounts for observable intermediate states (e.g., impregnating Cassie), and further identifies a new limiting state with potential realizability: a Cassie state accompanied by a precursor film, here termed the Inverse Wenzel state.
2 Theoretical Derivation
The derivation presented in this section is valid for predicting the local static contact angle when a volatile liquid wetting a solid surface is in chemical, thermal, and mechanical equilibrium with an ambient air/vapor phase. As illustrated in Fig. 1a, the liquid phase lies on the Bulk (B) side of the contact line and the vapor phase on the Ambient (A) side. We consider a chemically homogeneous substrate with (random or regular) three-dimensional micro- and nanoscale topographic features (Fig. 1a) producing a full surface area $S\geq A$, larger than the projected surface area $A$.
@@ 2 Theoretical Derivation
The full surface area is generally decomposed as $S=S_{\text{top}}+S_{\text{lat+bot}}$, where $S_{\text{top}}\!\geq 0$ is the planar area of feature tops, if present at all, and $S_{\text{lat+bot}}\!>0$ denotes the lateral-plus-basal area generated by the topography. We thus define the solid-top area fraction $\varphi_{S}=S_{\text{top}}/A$ and the lateral-basal area fraction $\Lambda=S_{\text{lat+bot}}/A$. With this notation, the Wenzel roughness ratio in Eq. 2 is $r=S/A=\varphi_{S}+\Lambda$. We will define $\varphi_{L}(h)=A_{\ell v}/A$ as the area fraction that would correspond to liquid-vapor interface if the surface topography were fully infiltrated by liquid up to the reference height $h$ of micro- or nanoscale dimensions. We will additionally assume that the reference liquid height $h$ within the infiltrated topography sets the local contact line position and is nearly constant within small distances $\lambda\sim h$ from the contact line. Note that, under assumptions leading to the Cassie-Baxter model (Eq. 3), we have $\varphi_{L}=1-\varphi_{S}$ but this equality is strictly valid when all topographic features have approximately the same height $h_{r}$ and the liquid-vapor interface is perfectly flat and coincides with the reference level $h=h_{r}$; see Ref. 49 for a detailed discussion. For random self-affine micro/nanoscale topography with tall peaks of small lateral extent protruding above the r.m.s. height, one may approximate $\varphi_{L}\simeq 1-\varphi_{S}$; an error on the order of roughly $15\%$ can be estimated for Gaussian height statistics.
Figure 1: Generalized model for static contact angles on a surface with micro/nanoscale surface topography. (a) General wetting state and unified model for varying degrees of liquid infiltration in the bulk side ($0\leq\phi_{B}\leq 1$) and ambient side ($0\leq\phi_{A}\leq 1$) of the contact line. (b) Wenzel state corresponding to full liquid infiltration under the liquid phase ($\phi_{B}=0$) and a perfectly dry surface under the ambient phase ($\phi_{A}=1$). (c) Cassie-Baxter state corresponding to strictly no liquid infiltration under the liquid and ambient phases ($\phi_{B}=\phi_{A}=1$). (d) Hemiwicking state corresponding to fill liquid infiltration under the liquid and ambient phases ($\phi_{B}=\phi_{A}=0$). (e) Inverse Wenzel state corresponding to strictly no liquid infiltration under the liquid phase ($\phi_{B}=1$) and full liquid infiltration under the ambient phase ($\phi_{A}=0$).
@@
Departing from the conventional treatment that leads to the classical Wenzel and Cassie-Baxter states, we consider that in general wetting and infiltration of the surface topography may not be ideally uniform; near the contact line, localized air/vapor voids may exist beneath the liquid phase, while localized liquid infiltration by wicking and condensation may occur through the texture on the ambient side exposed to the ambient vapor phase (see Fig. 1a). Hence, let $\phi_{B}\in[0,1]$ denote the averaged fraction of surface features not accessed by the liquid or partially occupied by air/vapor in the bulk side of the contact line and $\phi_{A}\in[0,1]$ denote the averaged fraction of surface features accessed and partially filled by liquid in the ambient side of the contact line. The limit condition with zero liquid-vapor interfacial area $S_{\ell v}=0$ on the bulk and ambient side thus correspond to $\phi_{R}=0$ and $\phi_{A}=1$; i.e., the Wenzel state (Fig. 1b). Similarly, the limit with $\phi_{R}=1$ and $\phi_{A}=1$ corresponds to the case that the vapor-liquid interface is $S_{\ell v}=\varphi_{L}A$; i.e., the Cassie-Baxter state (Fig. 1c).
Free-energy change and stationarity.
Following a conventional thermodynamic argument, we will consider that receding or advancing displacements of the contact line, under constant temperature, volume, and chemical equilibrium, produce a change in the free energy (or grand potential)42; 50 $dF=\gamma_{\ell v}\cos\theta+\gamma_{s\ell}dS_{s\ell}+\gamma_{sv}dS_{sv}+\gamma_{\ell v}dS_{\ell v}$, where $\gamma_{s\ell}$, $\gamma_{sv}$, and $\gamma_{\ell v}$ are the solid-liquid, solid-vapor, and liquid-vapor interfacial surface energies, and $S_{s\ell}$, $S_{sv}$, and $S_{\ell v}$ are the corresponding interfacial surface areas. Based on the geometric considerations above, a local displacement of the contact line over a projected area differential $\Delta A$, produces the following interfacial areas change:
@@
The general model for the local contact angle derived in Eq. 7 recovers the limiting behaviors predicted by previous well-established models, by making the conventional assumption that $\varphi_{L}+\varphi_{S}=1$: in the Wenzel state with $\phi_{B}=0$ and $\phi_{A}=1$, Eq. (7) reduces to Eq. 2 (Fig. 1b); in the Cassie-Baxter state for which $\phi_{B}=1$ and $\phi_{A}=1$ and the surface topography is fully filled by air/vapor on both the bulk and ambient sides, Eq. 7 yields Eq. 3 (Fig. 1c); and in the hemiwicking configuration (Fig. 1d), for which the surface topography is fully infiltrated by liquid on both sides adjacent to the contact line ($\phi_{B}=0$ and $\phi_{A}=0$), Eq. (7) simplifies to $\cos\theta=\varphi_{S}\cos\theta_{Y}+(1-\varphi_{S})$.42 In addition, Eq. 7 predicts a fourth limiting state, here termed the inverse Wenzel state (Fig. 1e), for which $\phi_{B}=1$ and $\phi_{A}=0$ and Eq. 7 gives
@@ 3 Results and Discussion
This section presents analytical results based on Eq. 7 for substrates with varying wettability, as characterized by the Young angle $\theta_{Y}$, and different surface topographies produced by random and engineered features. The analysis covers the full range of surface wetting conditions characterized by the liquid infiltration fractions on both the bulk and ambient sides adjacent to the contact line. For simplicity, the results reported here (Figs. 2–4) assume $\varphi_{L}=1-\varphi_{S}$. The implications of this approximation are discussed in the conclusions section.
@@
Figure 2: Sharp random topography. Local static contact angle $\theta$  predicted by Eq. 7 as a function of liquid-vapor surface fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surface topography with a small solid-top area fraction $\varphi_{S}=0.01$ and for two different lateral-plus-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
Sharp random topography.
@@
The results in Fig. 2 correspond to surfaces with sharp topographic features, with a small solid-top area fraction $\varphi_{S}=0.01$ and moderate-to-large lateral-plus-basal area fractions $\Lambda=1$ and $\Lambda=2$. The combination of topographic parameters in Fig. 2 represents, for example, common glass surfaces with natural micro/nanoscale roughness16; 51; 18 or engineered nanostructures with random nanomaterial deposition.52; 53 The analytical results in Fig. 2 show that a wide range of static contact angles, which deviate largely from the intrinsic Young contact angle, is feasible for this type of surface topography. For sufficiently “tall” topographic features giving large lateral area fractions $\Lambda\geq 2$, robust superhydrophilicity with $\theta\to 0$ can be realized for $\phi_{B}=0$ and $0\leq\phi_{A}\leq 1$ on wettable substrates with $\theta\lesssim 40^{\circ}$. This indicates that superwetting with vanishing hysteresis can be attained regardless of the conditions on the surface exposed to the ambient phase. Notably, large lateral area fractions $\Lambda\geq 2$ in a hydrophilic substrate can promote superhydrophobic states with $\theta\geq 150^{\circ}$ if the liquid phase is nearly fully suspended under topographically trapped air ($\phi_{B}\geq 0.9$), for arbitrary wetting conditions on the ambient side ($0\leq\phi_{A}\leq 1$). This surprising result correspond to the so-called inverse Wenzel state identified in Fig. 1e.
Engineered pillared structures.
@@
Figure 3: Engineered pillared structures. Local static contact angle $\theta$  predicted by Eq. 7 as a function of liquid-vapor surface fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surfaces with sharp topographic features, with a small solid-top area fraction $\varphi_{S}=0.5$ and for two different lateral-plus-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
@@
The static contact angle predictions in Fig. 3 correspond to surface topographies with a moderately large top-solid area fraction, $\varphi_{S}=0.5$, and lateral-plus-basal area fractions, $\Lambda=1$ and $\Lambda=2$. This type of surface topography is typically attained through engineered surface structures of micro- or nanoscale dimensions, featuring flat top surface areas with controlled dimensions and lateral-basal surface areas determined by the feature height and period (see, for example, Refs. 54; 9; 10). In this case, analytical predictions show less sensitivity to variations in the wetting conditions on the surface exposed to the liquid bulk and ambient vapor phase, resulting in relatively moderate hysteresis ranges for Young angles, $\theta_{Y}$, between 60 and 120, and shallow surface features for which $\Lambda\simeq 1$ (cf. Fig. 3). As shown in Fig. 3, surfaces with a moderately large top-solid area fraction, $\varphi_{S}=0.5$, require low Young angles, $\theta_{Y}\lesssim 40^{\circ}$, and tall structures with short periods that yield large lateral-basal area fractions, $\Lambda\gtrsim 2$, to produce robust superhydrophilicity for varying conditions on the surface side exposed to the ambient phase. The realizability of large contact angles, $\theta\geq 150^{\circ}$, for hydrophilic substrates with $\theta_{Y}\simeq 40^{\circ}$ is confined to a very narrow range of conditions ($\phi_{B}>0.9$ and $\phi_{A}<0.1$), for which the so-called inverse Wenzel state can exist. In addition, sticky superhydrophobicity with $\theta\geq 150^{\circ}$ can be attained on this type of surface topography (cf. Fig. 3) when the substrate is highly hydrophobic ($\theta_{Y}\geq 120^{\circ}$) and the liquid bulk infiltrates the topography underneath while the ambient side is partially wetted ($\phi_{B}\leq 0.5$ and $\phi_{A}\geq 0.9$).
Micro/nanocavity arrays.
The final cases analyzed (cf. Fig. 4) correspond to a large top-solid flat area fraction, $\varphi_{S}=0.8$, and lateral-plus-basal area fractions $\Lambda=1$ and $\Lambda=2$. This type of surface topography can be attained, for example, by patterning a smooth and flat surface with arrays of cavities with controllable micro- or nanoscale dimensions (e.g., Refs. 55; 56; 57). Note that if the cavities are not interconnected, capillary condensation and/or direct prewetting of the surface, rather than hemiwicking, are feasible mechanisms for infiltrating the surface topography. For this type of topography, robust superhydrophilicity with $\theta\to 0$ can be realized for $\phi_{A}\gtrsim 0.5$ and $\phi_{B}\lesssim 0.5$ in the case of large lateral-basal areas with $\Lambda>1$ and hydrophilic substrates with $\theta_{Y}\lesssim 40^{\circ}$. Superhydrophobicity is more difficult to achieve with this type of surface topography, and $\theta\geq 150^{\circ}$ is predicted only for highly hydrophobic substrates when $\phi_{A}\gtrsim 0.8$ and $\phi_{B}\lesssim 0.5$, which represent a rather narrow range of conditions.

### v2

@@ 1 Introduction
Prior work has focused on developing unified models that bridge the pure Wenzel and Cassie-Baxter extremes and extend these classical models.46; 47; 48 These efforts employ free energy minimization to account for multiple wetting states and assess their thermodynamic stability, for specific wetting configurations and surface geometries such as a macroscale sessile droplet on much smaller micropillar arrays. Despite the significant progress on the fundamental and applied aspects, open questions remain, namely how to predict static contact angles and the hysteresis range from a knowledge of topographic parameters, surface chemistry aging, and ambient conditions affecting the degree of liquid infiltration on the surface below and ahead of the wetting liquid phase. In an effort to address this, this work proposes a general and simple-to-use model, expanding Young’s local thermodynamic approach to consider surfaces with topographic features of general geometry and varying degrees of liquid infiltration. The introduced unified model recovers classical wetting limits (Wenzel, Cassie-Baxter, and hemiwicking states), accounts for observable intermediate states (e.g., impregnating Cassie), and further identifies a new limiting state with potential realizability: a Cassie state accompanied by a precursor film, here termed the Inverse Wenzel state.
The introduced model aims to predict the static contact angle $\theta$ from the substrate and interfaces chemistry, characterized by the Young contact angle $\theta_{Y}$, and geometric parameters of the surface topography, such as the top, lateral, and basal surface areas, for generalized wetting conditions prescribed by arbitrary degrees of liquid infiltration beneath the bulk liquid and ambient phases. It is worth noting that the liquid infiltration fractions on the surface topography adjacent to the contact line, treated as known inputs in this derivation, result from kinetically trapped (metastable) or thermodynamically stable states that can coexist. These infiltration fractions may or may not be realized under the specific physicochemical conditions and/or dynamic constraints in the actual system studied, as in the case of the classical Wenzel and Cassie-Baxter models (Eqs. 23).
As illustrated in Fig. 1a, the liquid phase lies on the Bulk (B) side of the contact line and the vapor phase on the Ambient (A) side. We consider a chemically homogeneous substrate with (random or regular) three-dimensional micro- and nanoscale topographic features (Fig. 1a) producing a full surface area $S\geq A$, larger than the projected surface area $A$.
@@
The full surface area is generally decomposed as $S=S_{\text{top}}+S_{\text{lat+bot}}$, where $S_{\text{top}}\!\geq 0$ is the planar area of feature tops, if present at all, and $S_{\text{lat+bot}}\!>0$ denotes the lateral plus basal area generated by the topography. We thus define the solid-top flat area fraction $\varphi_{S}=S_{\text{top}}/A$ and the lateral-basal area fraction $\Lambda=S_{\text{lat+bot}}/A$. With this notation, the Wenzel roughness ratio in Eq. 2 is $r=S/A=\varphi_{S}+\Lambda$. We will define $\varphi_{L}(h)=A_{\ell v}/A$ as the area fraction that would correspond to liquid-vapor interface if the surface topography were fully infiltrated by liquid up to the reference height $h$ of micro- or nanoscale dimensions. We will additionally assume that the reference liquid height $h$ within the infiltrated topography sets the local contact line position and is nearly constant within small distances $\lambda\sim h$ from the contact line. Note that, under assumptions leading to the Cassie-Baxter model (Eq. 3), we have $\varphi_{L}=1-\varphi_{S}$ but this equality is strictly valid when all topographic features have approximately the same height $h_{r}$ and the liquid-vapor interface is perfectly flat and coincides with the reference level $h=h_{r}$; see Ref. 49 for a detailed discussion. For random self-affine micro/nanoscale topography with tall peaks of small lateral extent protruding above the r.m.s. height, one may approximate $\varphi_{L}\simeq 1-\varphi_{S}$; an error on the order of roughly $15\%$ can be estimated for Gaussian height statistics.
Figure 1: Unified model for static contact angles on a surface with micro/nanoscale surface topography. (a) Generalized wetting state and unified model for arbitrary degrees of liquid infiltration in the bulk side ($0\leq\phi_{B}\leq 1$) and ambient side ($0\leq\phi_{A}\leq 1$) of the contact line. (b) Wenzel state corresponding to full liquid infiltration under the liquid phase ($\phi_{B}=0$) and a perfectly dry surface under the ambient phase ($\phi_{A}=1$). (c) Cassie-Baxter state corresponding to strictly no liquid infiltration under the liquid and ambient phases ($\phi_{B}=\phi_{A}=1$). (d) Hemiwicking state corresponding to full liquid infiltration under the liquid and ambient phases ($\phi_{B}=\phi_{A}=0$). (e) Inverse Wenzel state corresponding to strictly no liquid infiltration under the liquid phase ($\phi_{B}=1$) and full liquid infiltration under the ambient phase ($\phi_{A}=0$).
@@
Deviating from the conventional treatment that leads to the classical Wenzel and Cassie-Baxter states, we consider that in general wetting and infiltration of the surface topography may be partial and nonuniform; near the contact line, localized air/vapor voids may exist beneath the liquid phase, while localized liquid infiltration by wicking and condensation may occur through the texture on the ambient side exposed to the ambient vapor phase (see Fig. 1a). Hence, let $\phi_{B}\in[0,1]$ denote the averaged fraction of surface features occupied by air/vapor in the bulk side of the contact line and $\phi_{A}\in[0,1]$ denote the averaged fraction of surface features filled with air/vapor in the ambient side of the contact line. The limit condition with zero liquid-vapor interfacial area $S_{\ell v}=0$ on the bulk and ambient side thus correspond to $\phi_{B}=0$ and $\phi_{A}=1$; i.e., the Wenzel state (Fig. 1b). Similarly, the limit with $\phi_{B}=1$ and $\phi_{A}=1$ corresponds to the case that the liquid-vapor interface is $S_{\ell v}=\varphi_{L}A$; i.e., the Cassie-Baxter state (Fig. 1c).
Free-energy stationarity.
Following a conventional thermodynamic argument, we will consider that receding or advancing displacements of the contact line, under constant temperature, volume, and chemical equilibrium, produce a change in the free energy (or grand potential)42; 50 $dF=\gamma_{\ell v}\cos\theta dA+\gamma_{s\ell}dS_{s\ell}+\gamma_{sv}dS_{sv}+\gamma_{\ell v}dS_{\ell v}$, where $\gamma_{s\ell}$, $\gamma_{sv}$, and $\gamma_{\ell v}$ are the solid-liquid, solid-vapor, and liquid-vapor interfacial surface energies, and $S_{s\ell}$, $S_{sv}$, and $S_{\ell v}$ are the corresponding interfacial surface areas. Based on the geometric considerations above, a local displacement of the contact line over a projected area differential $\Delta A$, produces the following interfacial areas change:
@@
The general model for the local contact angle derived in Eq. 7 recovers the limiting behaviors predicted by previous well-established models, by making the conventional assumption that $\varphi_{L}+\varphi_{S}=1$: in the Wenzel state with $\phi_{B}=0$ and $\phi_{A}=1$, Eq. 7 reduces to Eq. 2 (Fig. 1b); in the Cassie-Baxter state for which $\phi_{B}=1$ and $\phi_{A}=1$ and the surface topography is fully filled with air/vapor on both the bulk and ambient sides, Eq. 7 yields Eq. 3 (Fig. 1c); and in the hemiwicking configuration (Fig. 1d), for which the surface topography is fully infiltrated by liquid on both sides adjacent to the contact line ($\phi_{B}=0$ and $\phi_{A}=0$), Eq. (7) simplifies to $\cos\theta=\varphi_{S}\cos\theta_{Y}+(1-\varphi_{S})$.42 In addition, Eq. 7 predicts a fourth limiting state, here termed the Inverse Wenzel state (Fig. 1e), for which $\phi_{B}=1$ and $\phi_{A}=0$ and Eq. 7 gives
@@ 2 Results and Discussion
This section presents analytical results from Eq. 7 for a wide range of physically meaningful Young contact angles $\theta_{Y}$ and different surface topographies characterized by their lateral-basal area fraction $\Lambda$ and top-solid flat area fraction $\varphi_{S}$. The analysis covers the full range of topography infiltration conditions, characterized by the liquid-vapor area fractions on the bulk and ambient sides adjacent to the contact line. For simplicity, the results reported here assume $\varphi_{L}\simeq 1-\varphi_{S}$.
2.1 Sharp random topography
@@
Figure 2: Sharp random topography. Local static contact angle $\theta$  predicted by Eq. 7 as a function of vapor fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surface topography with a small solid-top flat area fraction $\varphi_{S}=0.01$ and for two different lateral-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
@@
The results in Fig. 2 correspond to surfaces with sharp topographic features, with a small solid-top flat area fraction $\varphi_{S}=0.01$ and moderate-to-large lateral-basal area fractions $\Lambda=1$ and $\Lambda=2$. The combination of topographic parameters in Fig. 2 represents, for example, common glass surfaces with natural micro/nanoscale roughness16; 51; 18 or engineered nanostructures with random nanomaterial deposition.52; 53 The analytical results in Fig. 2 show that a wide range of static contact angles, which deviate largely from the intrinsic Young contact angle, is feasible for this type of surface topography. For sufficiently “tall” topographic features giving large lateral-basal area fractions $\Lambda\geq 2$, robust superhydrophilicity with $\theta\to 0$ can be realized for $\phi_{B}=0$ and $0\leq\phi_{A}\leq 1$ on wettable substrates with $\theta\lesssim 40^{\circ}$. This indicates that superwetting with vanishing hysteresis can be attained regardless of the conditions on the surface exposed to the ambient phase. Notably, large lateral-basal area fractions $\Lambda\geq 2$ in a hydrophilic substrate can promote superhydrophobic states with $\theta\geq 150^{\circ}$ if the liquid phase is nearly fully suspended under topographically trapped air ($\phi_{B}\geq 0.9$), for arbitrary wetting conditions on the ambient side ($0\leq\phi_{A}\leq 1$). This surprising result corresponds to the so-called Inverse Wenzel state identified in Fig. 1e.
2.2 Engineered pillared structures
@@
Figure 3: Engineered pillared structures. Local static contact angle $\theta$  predicted by Eq. 7 as a function of vapor fractions $\phi_{B}$ (bulk) and $\phi_{A}$ (ambient) adjacent to the contact line, for three Young angles: $\theta_{Y}=40^{\circ},\,90^{\circ},\,\&\,120^{\circ}$. The results correspond to surfaces with sharp topographic features, with a small solid-top flat area fraction $\varphi_{S}=0.5$ and for two different lateral-basal area fractions $\Lambda=1$ (top row) and $\Lambda=2$ (bottom row). Wetting states: (W) Wenzel, (CB) Cassie-Baxter, (IW) Inverse Wenzel, and (H) Hemiwicked.
@@
The static contact angle predictions in Fig. 3 correspond to surface topographies with a moderately large top-solid area fraction, $\varphi_{S}=0.5$, and lateral-basal area fractions, $\Lambda=1$ and $\Lambda=2$. This type of surface topography is typically attained through engineered surface structures of micro- or nanoscale dimensions, featuring flat top surface areas with controlled dimensions and lateral-basal surface areas determined by the feature height and period (see, for example, Refs. 54; 9; 10). In this case, analytical predictions show less sensitivity to variations in the wetting conditions on the surface exposed to the liquid bulk and ambient vapor phase, resulting in relatively moderate hysteresis ranges for Young angles, $\theta_{Y}$, between 60 and 120, and shallow surface features for which $\Lambda\simeq 1$ (cf. Fig. 3). As shown in Fig. 3, surfaces with a moderately large top-solid area fraction, $\varphi_{S}=0.5$, require low Young angles, $\theta_{Y}\lesssim 40^{\circ}$, and tall structures with short periods that yield large lateral-basal area fractions, $\Lambda\gtrsim 2$, to produce robust superhydrophilicity for varying conditions on the surface side exposed to the ambient phase. The realizability of large contact angles, $\theta\geq 150^{\circ}$, for hydrophilic substrates with $\theta_{Y}\simeq 40^{\circ}$ is confined to a very narrow range of conditions ($\phi_{B}>0.9$ and $\phi_{A}<0.1$), for which the so-called Inverse Wenzel state can exist. In addition, sticky superhydrophobicity with $\theta\geq 150^{\circ}$ can be attained on this type of surface topography (cf. Fig. 3) when the substrate is highly hydrophobic ($\theta_{Y}\geq 120^{\circ}$) and the liquid bulk infiltrates the topography underneath while the ambient side is partially wetted ($\phi_{B}\leq 0.5$ and $\phi_{A}\geq 0.9$).
2.3 Micro/nanocavity arrays